# L-MARS: Legal Multi-Agent Workflow with Orchestrated Reasoning and Agentic Search

### v2

@@ 4.3 Implementation Details
All systems were evaluated on the 200-question LegalSearchQA benchmark using a standardized pipeline. The baseline LLMs (GPT-4o, Claude-4-Sonnet, Gemini-2.5-Flash) were run with temperature =0=0 and a 2048 token limit, using structured output or JSON mode to ensure consistency. Each model produced both a multiple choice answer and a supporting explanation, which were aggregated into structured JSON for evaluation. Performance was assessed along three complementary axes: accuracy, U-Score, and LLM-as-Judge. For reporting, accuracy and U-Score are averaged across all questions, while LLM-as-Judge ratings are determined by majority vote among per-question judgments. The complete evaluation pipeline, including prompts and scripts, is provided in the eval/ directory of our repository.
4.4 Results
Main Results on LegalSearchQA
Table 3 reports aggregate results across three conditions: pure LLM inference, L-MARS in simple mode, and L-MARS in multi-turn mode. Our results highlight three key findings. First, both L-MARS variants substantially outperform pure LLM inference in accuracy (up to 98% vs. 86–89%). Second, the U-Score decreases sharply under L-MARS (from 0.55–0.62 to 0.39–0.42), showing a marked reduction in hedging, vagueness, and unsupported conclusions. Third, LLM-as-Judge ratings consistently rank L-MARS outputs higher, with the multi-turn variant producing more thorough and contextually grounded answers. Finally, we note a trade-off in response time: while baseline LLMs answer within 1-4 seconds, L-MARS incurs higher latency (13.6s for simple, 55.7s for multi-turn), reflecting the added retrieval and reasoning steps. The visualization of the results is shown in Figure 3, and a detailed case study can be found in Appendix B.
Table 3: Average performance of baseline LLMs and our proposed L-MARS framework on the LegalSearchQA benchmark.
Method Accuracy U-Score LLM-as-Judge Response Time (s)
GPT-4o 0.89 0.55 Moderate 1.69
Claude-4-Sonnet 0.88 0.62 Moderate 3.84
Gemini-2.5-Flash 0.86 0.58 Moderate 1.87
L-MARS (simple, ours) 0.96 0.42 High 13.62
L-MARS (multi-turn, ours) 0.98 0.39 High 55.67
@@ 4.5 Error Analysis
To better understand the limitations of the model, we conducted a qualitative error analysis of incorrect predictions. We observed two recurring patterns. First, models sometimes default to heuristic reasoning, such as assuming generic timelines or standard legal procedures rather than following the specific statutory text. Second, models occasionally overgeneralized from frequent legal phrases, leading to misclassification when a question involved narrow carving-outs or exceptions. These errors highlight two key challenges: reliance on surface-level priors instead of precise retrieval, and difficulty distinguishing closely related legal categories. Both issues underscore the importance of retrieval grounding and calibrated reasoning in complex legal QA.
Figure 3: Average performance of baseline LLMs and our proposed L-MARS framework on the LegalSearchQA benchmark. Bars report Accuracy (blue) and $1-\text{U-Score}$ (purple), with response times shown alongside the $y$-axis labels. L-MARS substantially improves factual correctness and reduces uncertainty, though with higher latency compared to pure LLM inference.
Human Expert Evaluation
In addition, we conducted blind evaluations with graduate law students on a subset of 50 questions. Human preferences were closely aligned with LLM-as-Judge outcomes, with a 0.92 inter-annotator agreement rate. A recurring failure case was the Judge Agent’s tendency to over-reject partially relevant sources, occasionally triggering unnecessary search iterations. This suggests a direction for future work on trajectory-aware judge models that retain awareness of prior retrieval steps.
@@ 5 Conclusion
We introduced L-MARS, a multi-agent workflow that integrates structured reasoning, agentic search, and sufficiency verification for legal question answering. By orchestrating specialized agents in iterative loops of query refinement, targeted retrieval, and evidence-based judgment, L-MARS significantly reduces hallucinations and uncertainty compared to pure LLM baselines. Our evaluation of the LegalSearchQA benchmark demonstrates consistent gains in factual accuracy, decisiveness, and grounding in authoritative sources.
6 Limitations and Future Work
While L-MARS improves reliability, its performance remains bounded by retrieval quality: if search engines or legal databases fail to return relevant authorities, downstream reasoning is impaired. The iterative multi-turn mode also introduces higher latency, which may be prohibitive in real-time legal assistance settings. Moreover, our evaluation focuses on US law; broader cross-jurisdictional and multilingual evaluations are needed to assess generalizability. Finally, the Judge Agent, though effective at flagging insufficiency, can be overly conservative, sometimes triggering unnecessary search iterations.
@@ Appendix A Agent Instructions
Query Agent Prompts
# Follow-up Questions Generation Prompt
You are a legal assistant helping users with their legal questions.
User’s original question: {user_query}
Generate 2-3 clarifying follow-up questions that would help you provide better
legal assistance.
Focus on:
- Jurisdiction/location if not specified
- Specific circumstances or context
- Timeline or urgency
- What type of help they need (information, next steps, etc.)
Only ask questions that would significantly help in providing better assistance.
# Search Query Generation Prompt
Based on the user’s legal question and any additional context, generate 2-4 specific
search queries.
User Question: {user_query}
Additional Context: {context}
Generate queries for different sources:
- case_law: For searching legal cases and precedents
- web_search: For general legal information and recent updates
- offline_rag: For searching local legal documents in the inputs folder
Make queries specific and focused to get the most relevant results.
# Query Analysis Prompt
Analyze the user’s legal query and provide a detailed structured breakdown.
User Query: {user_query}
Additional Context: {context}
Analyze and extract:
1. Legal category/area (e.g., contract law, criminal law, family law, employment law)
2. People involved with their roles (plaintiff, defendant, client, witness, etc.)
3. Jurisdiction if mentioned or inferable
4. Urgency level based on the nature of the query
5. Specific legal areas involved
6. Timeline or deadlines if mentioned
7. Additional context about the situation
Be thorough but accurate in your categorization.
Summary Agent Prompt
Create a comprehensive answer to the user’s legal question based on the search results.
User’s Question: {user_query}
Search Results:
{results_content}
Provide:
1. A clear, comprehensive answer
2. Key legal points and considerations
3. Important disclaimers about legal advice
Remember: This is informational only and not legal advice.
Judge Agent Prompt
You are a legal research judge evaluating search results.
This is iteration {iteration_count + 1}.
Original Question: {user_query}
Conversation History:
{conversation_context}
Current Search Results ({len(search_results)} results):
{results_summary}
{prev_evaluations}
EVALUATION PROTOCOL WITH CHAIN-OF-THOUGHT:
1. REASONING (Chain of Thought):
Think step by step about whether the search results answer the user’s specific question.
Consider: What was asked? What information was provided? What is still missing?
2. SOURCE QUALITY CHECK:
Analyze source authority - are there sources from:
- Government (.gov) sites?
- Court decisions/legal databases?
- Educational institutions (.edu)?
- How many authoritative vs user-generated content sources?
3. DATE CHECK:
- Are the sources current and relevant to today’s date?
- If there are older sources, do we also have recent confirmations?
- Flag if critical information might be outdated
4. JURISDICTION CHECK:
- Does the jurisdiction of sources match the user’s location/scope?
- For US federal vs state law, is the distinction clear?
- User mentioned: {conversation_context if conversation_context
else "No specific jurisdiction mentioned"}
5. CONTRADICTION SCAN:
- Do any sources contradict each other?
- If yes, what specific elements conflict?
- Do we need more specific queries to resolve conflicts?
STOP RULE:
Mark as SUFFICIENT when:
- Primary claims have authoritative support (gov/court/edu when available)
- No critical information gaps for practical guidance
- No unresolved contradictions
- Jurisdiction and dates are appropriate
Mark as INSUFFICIENT when:
- Missing critical legal requirements or procedures
- Only user-generated content for key claims
- Significant contradictions need resolution
- Wrong jurisdiction or outdated information
IMPORTANT for iteration {iteration_count + 1}:
- Be MORE LENIENT after multiple iterations
- Focus on whether user has enough info to take action
- Consider cumulative information across all iterations
SUGGESTED REFINEMENTS (if insufficient):
Provide 1-2 specific search queries that would help find the missing information.
These should be actual search queries that can be executed directly, NOT instructions.
Good examples:
- "California SB 365 employment arbitration 2025 full text"
- "Biden student loan forgiveness Supreme Court ruling 2024"
- "AB 465 arbitration agreement requirements California"
Bad examples (don’t do this):
- "Search for more information about SB 365"
- "Look for court rulings"
- "Find government sources"
Provide detailed reasoning for your judgment.
Appendix B Case Study: Executive Order Timeline Interpretation
This section presents a comparative analysis of a legal multiple-choice question where the L-MARS system with retrieval-augmented generation successfully identified the correct answer, while GPT-4o without external search failed.
Question ID 32: Executive Order Timeline
Question: According to Section 3 of the May 23, 2025 Executive Order
’Restoring Gold Standard Science,’ how soon must the OSTP Director
issue guidance for agencies on implementing ’Gold Standard Science’?
Answer Choices:
A: Within 30 days of the date of the order
B: Within 90 days of the date of the order
C: Within 6 months of the date of the order
D: Within 1 year of the date of the order
Correct Answer: A
L-MARS Simple Mode Response (Correct)
Selected Answer: A (Within 30 days)
Reasoning: The search results from library.washu.edu explicitly state
that the order mandated a timeline of 30 days for the Office of Science
and Technology Policy to issue guidance. This is the most direct evidence
relating to Section 3 of the Executive Order. The system cross-referenced
multiple authoritative sources to confirm this specific timeline requirement
rather than relying on general patterns or assumptions about typical
executive order timelines.
Search Sources Used: 5 Web Search sources
1. whitehouse.gov - "Restoring Gold Standard Science"
   Official White House publication of the Executive Order with full text
2. justice.gov - "New Executive Order on ’Gold Standard Science’: FOIA..."
   Department of Justice guidance on FOIA implications and compliance timelines
3. $h$s.gov - "Implementing Gold Standard Science"
   Department of Health and Human Services implementation guidelines
4. lawbc.com - "White House OSTP Issues Agency Guidance for Gold..."
   Legal analysis of OSTP guidance requirements and deadlines
5. library.washu.edu - "Federal Agencies Respond to ’Gold Standard Science..."
   Academic repository with detailed timeline analysis stating explicitly:
   "Section 3 mandates 30-day timeline for OSTP guidance issuance"
Key Evidence Retrieved:
- Executive Order issued: May 23, 2025
- OSTP guidance deadline: Within 30 days (by June 22, 2025)
- Agencies must implement principles of scientific integrity
- Guidance covers conflicts of interest in research and policy
Performance Metrics:
- Processing Time: 13.08 seconds
- Confidence Score: 0.9 (90%)
- U-Score (Uncertainty): 0.611 (lower is better)
- Citation Score: 1.0 (perfect source attribution)
- Jurisdiction Score: 1.0 (correctly identified federal scope)
- Decisiveness Score: 0.5 (moderate - provided clear guidance)
- Model: OpenAI GPT-4o (with RAG)
Result: Correct
GPT-4o Response (Incorrect)
Selected Answer: B (Within 90 days)
Reasoning: The question specifies the timing requirement for issuing
guidance by the OSTP Director according to Section 3 of the Executive
Order. Based on the context and typical executive orders that set
implementation timelines for guidance issuance, 90 days is commonly
allocated as a reasonable period for action. Therefore, the best
answer is B: Within 90 days of the date of the order.
Result: Incorrect
@@ B.1 Analysis
This case demonstrates a critical advantage of retrieval-augmented generation in legal question answering:
B.1.1 Why GPT-4o Failed
  1. 1.
    Reliance on General Patterns: GPT-4o relied on typical patterns in executive orders, assuming 90 days as a “reasonable period” without access to the specific order text.
  2. 2.
    Lack of Specific Knowledge: The model’s training data cutoff prevented it from having specific information about an executive order in May 2025.
  3. 3.
    Statistical Bias: The model defaulted to a common timeline (90 days) that appears frequently in government documents, demonstrating the limitation of pattern-based reasoning without factual verification.
B.1.2 Why L-MARS Succeeded
  1. 1.
    Direct Evidence Retrieval: L-MARS retrieved the actual executive order text from authoritative sources, finding the explicit 30-day requirement in Section 3.
  2. 2.
    Source Authority: The system identified and prioritized information from library.washu.edu, an authoritative academic source with direct access to government documents.
  3. 3.
    Evidence-Based Reasoning: Rather than relying on typical patterns, L-MARS based its answer on explicit textual evidence from the retrieved documents.
B.2 Implications for Legal AI Systems
This case study highlights several important considerations for legal AI applications.
  • Temporal Sensitivity: Legal questions often require access to current and specific documents that may not be found in training data.
  • Precision Requirements: Legal deadlines and requirements demand exact information rather than reasonable approximations.
  • Source Verification: The ability to cite specific authoritative sources is crucial for the credibility of legal research.
  • RAG Advantage: Retrieval-augmented generation provides a significant advantage for questions requiring specific factual information, particularly for recent or specialized legal documents.
This example underscores that while large language models possess strong reasoning capabilities, augmenting them with retrieval mechanisms is essential for reliable legal question answering, especially when dealing with specific statutory requirements, deadlines, and recent legal developments.

### v3

@@ 4.3 Implementation Details
All conditions were evaluated using a standardized pipeline. For LegalSearchQA, we evaluated all 50 questions under each of the three conditions. For Bar Exam QA, we evaluated the full 594-question subset under zero-shot and L-MARS conditions. For the zero-shot and CoT conditions, prompts instruct the model to select a single answer letter (A–D). For L-MARS, all agents use GPT-4o-mini with T=0T{=}0; CourtListener integration was disabled since neither benchmark requires specific case citations. On Bar Exam QA, we additionally tested L-MARS with the Tavily search backend as an alternative to Serper. The complete evaluation pipeline, including prompts, scripts, and all results, is available at https://github.com/boqiny/L-MARS.
5 Results
5.1 LegalSearchQA
Table 2 reports accuracy across three conditions on LegalSearchQA. L-MARS achieves 96.0% accuracy, improving over zero-shot by 38 percentage points while maintaining moderate latency at 2.4s per question. Chain-of-thought prompting degrades performance from 58.0% to 30.0%, falling below the 25% random baseline for four-choice questions. Without access to current information, CoT causes the model to construct plausible but outdated rationales and arrive at wrong answers with high confidence. We term this confident confabulation: the model recalls stale figures from prior tax years or superseded regulations, then builds step-by-step justifications around them. L-MARS rescued 19 questions that zero-shot answered incorrectly while regressing on none. We use a standard legal reasoning CoT prompt following the issue-rule-application-conclusion (IRAC) structure. We attribute the degradation to reliance on outdated parametric knowledge rather than prompt design, as the model produces internally coherent reasoning from stale premises across all five domains. We verified consistent degradation patterns with temperature T=0T{=}0 across all five domains, suggesting the effect is robust rather than an artifact of a single prompt configuration.
Table 2: Accuracy on LegalSearchQA (50 questions). All conditions use GPT-4o-mini.
Condition Retrieval Accuracy (%) Correct/Total Avg. Latency (s)
Zero-Shot None 58.0 29/50 0.7
Chain-of-Thought None 30.0 15/50 6.8
L-MARS (Simple) Serper + agents 96.0 48/50 2.4
Figure 2 and Table 3 break down accuracy by legal domain. The largest gains appear in Tax, Corporate & Financial Regulation, where L-MARS improves over zero-shot by 61.5 percentage points. Questions in this domain ask for specific dollar amounts, filing deadlines, and regulatory thresholds that changed in 2025 and 2026. Zero-shot accuracy here is only 30.8%, reflecting the model’s inability to access post-training information; L-MARS reaches 92.3% by retrieving current IRS and SEC sources. The smallest gains appear in Criminal, Drug & State Law, where zero-shot already achieves 87.5% because many questions concern well-publicized events that likely entered the model’s training data through news coverage. Notably, CoT falls below the random baseline in three of five domains, confirming that reasoning from stale knowledge is worse than no reasoning at all when current facts are required.
Figure 2: Per-domain accuracy on LegalSearchQA. L-MARS achieves near-perfect accuracy across all domains. CoT falls below the 25% random baseline in three of five domains.
Table 3: Per-domain accuracy (%) on LegalSearchQA. $\Delta$ is L-MARS gain over zero-shot.
Domain N Zero-Shot CoT L-MARS $\Delta$
Tax, Corporate & Financial 13 30.8 15.4 92.3 +61.5
Labor & Employment 11 63.6 18.2 100.0 +36.4
Technology & Privacy 9 55.6 44.4 100.0 +44.4
Immigration 9 66.7 22.2 88.9 +22.2
Criminal, Drug & State Law 8 87.5 62.5 100.0 +12.5
Overall 50 58.0 30.0 96.0 +38.0
5.2 Bar Exam QA
Figure 3 visualizes the contrast between the two benchmarks. On Bar Exam QA (Table 4), L-MARS provides a negligible 0.7 point gain with Serper and a slight 0.8 point degradation with Tavily. Bar Exam QA tests legal reasoning over hypothetical fact patterns whose answers follow from general legal principles already in the model’s parameters, so web retrieval adds noise rather than signal. The gap between a 38 point gain on LegalSearchQA and a 0.7 point gain on Bar Exam QA confirms that the value of agentic retrieval depends on whether the task requires current information or abstract reasoning.
Figure 3: Benchmark comparison. On retrieval-dependent LegalSearchQA, L-MARS gains 38 percentage points over zero-shot. On reasoning-focused Bar Exam QA, retrieval provides negligible improvement.
Table 4: Accuracy on Bar Exam QA (594 questions). All conditions use GPT-4o-mini.
Condition Acc. (%) Correct
Zero-Shot 55.2 328/594
L-MARS (Serper) 55.9 332/594
L-MARS (Tavily) 54.4 323/594
Prior work [28]:
GPT-4o-mini 49.7
+ Best retriever 50.2
Table 5: Per-subject accuracy (%) on Bar Exam QA.
Subject N ZS L-MARS
Const. Law 95 81.1 76.8
Contracts 113 46.9 52.2
Criminal 89 43.8 41.6
Evidence 93 54.8 61.3
Real Prop. 92 47.8 50.0
Torts 112 57.1 53.6
Table 5 shows per-subject results. L-MARS with Serper yields modest gains on Contracts (+5.3 points) and Evidence (+6.5 points) but slight regressions on Constitutional Law and Torts. No subject shows a large or consistent benefit from retrieval, reinforcing the conclusion that this benchmark primarily tests reasoning ability rather than information access.
@@ 5.3 Error Analysis
L-MARS answered only two LegalSearchQA questions incorrectly. Q27 asked about H-1B weighted selection details involving a nuanced procedural rule about wage-level assignment for multiple registrations; the retrieved results did not surface the specific Federal Register provision. Q44 asked for the exact effective date of T+1 settlement; search results discussed the policy broadly but missed the precise date. Both errors share a pattern: the correct answer requires a specific fact buried in a longer document, and snippet-level retrieval missed the critical sentence.
The CoT degradation deserves separate attention. On tax threshold questions, for instance, the model recalls outdated dollar amounts from prior years and builds multi-step justifications around them. This pattern is not unique to legal QA. In any domain where facts change faster than model training cycles, prompting the model to reason more carefully can amplify rather than correct errors, because the model reasons from wrong premises with greater conviction. This finding suggests caution when deploying CoT in fast-changing domains without concurrent retrieval.
5.4 Ablation: Retrieval Depth
[…]
@@ 6 Conclusion
We presented L-MARS, a multi-agent retrieval framework for legal QA, and LegalSearchQA, a 50-question benchmark requiring current legal knowledge that post-dates model training data. L-MARS achieves 96.0% on LegalSearchQA, a 38 percentage point gain over zero-shot, while CoT degrades accuracy to 30.0% through confident confabulation from stale knowledge. On reasoning-focused Bar Exam QA, retrieval provides negligible gains, confirming that the value of agentic retrieval is benchmark-dependent. These findings underscore the need for retrieval-focused legal benchmarks and suggest caution when deploying CoT in fast-changing domains without concurrent retrieval.
Limitations
LegalSearchQA contains 50 questions verified as of March 2026; the benchmark is sufficient for demonstrating large effect sizes but limits per-domain statistical power, and answers may become outdated as laws change. We evaluate only Simple Mode with web search; the Multi-Turn Mode, local RAG indexing, and CourtListener integration are available in the released codebase but not benchmarked here, as the simple configuration already demonstrates the core finding. Our evaluation covers U.S. law in English with a single backbone model, and performance remains bounded by search engine quality, as seen in the two errors where snippet-level retrieval missed critical details buried in longer documents.
@@ Appendix A Agent Instructions
Query Agent
The Query Agent is a stateless pass-through that wraps the user
query into a structured QueryGeneration dataclass:
    @dataclass
    class QueryGeneration:
        query: str
        query_type: str = "web_search"
        priority: str = "high"
    class QueryAgent:
        def build_query(self, user_query: str) -> QueryGeneration:
            return QueryGeneration(query=user_query.strip())
Summary Agent
You are a bar exam expert. A user will provide retrieved legal
sources and a multiple-choice question. Use the retrieved sources
together with your legal knowledge to select the single best answer.
Rules:
1. Your response MUST start with exactly one letter on the first
   line: A, B, C, or D.
2. Follow with a concise 1-2 sentence rationale citing the
   controlling legal rule.
3. Bar exam questions test specific exceptions and nuances -- do
   not just apply the most common general rule if the specific
   facts call for an exception.
Response format:
[Letter]
[Brief rationale]
The user message is structured as:
  Retrieved Evidence:
  <concatenated search results with source, title, content>
  Question:
  <full MCQ with answer choices A-D>
Judge Agent System Prompt
You are a legal research quality judge for US bar exam questions.
You will receive a bar exam multiple-choice question and a list of
retrieved search result snippets. Your job is to decide whether the
retrieved evidence is sufficient to reliably answer the question.
"Sufficient" means: at least one result contains specific legal
rules, statutes, case holdings, or doctrinal exceptions that
directly address the core legal issue tested by the question -- not
just a restatement of the facts.
If the evidence is NOT sufficient, provide a better search query
that targets primary legal sources (e.g., specific rule names,
statute citations, or doctrine names) rather than generic question
text.
Respond with valid JSON only, no markdown fences:
{"sufficient": true_or_false, "reason": "one sentence",
 "next_query": "refined query (omit if sufficient)"}
Pre-filtering: Before LLM evaluation, results matching any of
these patterns are deterministically removed at zero cost:
  HTTP error, 403/404, Connection error, Request timed out,
  JavaScript is disabled, No search results found, etc.
Appendix B Case Study: Post-Training Knowledge Gap
We present two representative examples from LegalSearchQA where zero-shot and chain-of-thought both fail due to outdated parametric knowledge, while L-MARS retrieves the correct current information.
Q1 (lsqa_001): Executive Order 14110 Status
Question: What is the current status of Executive Order 14110, the
Biden administration’s executive order on AI safety signed in
October 2023?
A. It remains in full effect and is being actively enforced
B. It was revoked by President Trump on January 20, 2025
C. It was partially amended but key provisions remain
D. It expired automatically after 12 months
Correct Answer: B
Zero-Shot Response (Incorrect)
Selected: A
The model’s parametric knowledge contains EO 14110 as an active,
enforceable order -- it has no information about the January 2025
revocation, which post-dates its training data.
Chain-of-Thought Response (Incorrect)
Selected: A
The model reasons step-by-step from stale knowledge: it recalls
that EO 14110 was signed in October 2023, notes that executive
orders do not expire automatically, and concludes it must still be
in effect. The reasoning is internally consistent but factually
wrong -- an example of "confident confabulation" from outdated
parametric memory.
L-MARS Response (Correct)
Selected: B
Executive Order 14110 has been revoked, as indicated by the sources
stating that President Trump took action to revoke it during his
final week in office. Therefore, it is no longer in effect.
Retrieved sources included whitehouse.gov and Federal Register
entries confirming revocation via EO 14148 on January 20, 2025.
Q2 (lsqa_002): 2025 Standard Deduction
Question: For tax year 2025, what is the standard deduction for a
single filer under age 65, after the One Big Beautiful Bill Act
amendments?
A. \$14,600    B. \$15,000    C. \$15,350    D. \$15,750
Correct Answer: D
Zero-Shot: C (Incorrect)   CoT: C (Incorrect)   L-MARS: D (Correct)
Zero-Shot selected C (\$15,350) -- the model recalls an intermediate
figure from pre-OBBBA IRS projections but lacks knowledge of the
legislative amendment.
CoT selected C (\$15,350) -- step-by-step reasoning reinforces the
wrong figure: the model recalls the 2024 standard deduction
(\$14,600), applies an estimated inflation adjustment, and arrives
at \$15,350. The reasoning is plausible but based on stale data.
L-MARS selected D (\$15,750) -- retrieved current IRS sources
confirming the One Big Beautiful Bill Act raised the 2025 standard
deduction to \$15,750 for single filers.
@@ B.1 Analysis
Both examples illustrate the same failure mode: the model’s parametric knowledge is confidently wrong about post-training developments. In Q1, a major policy reversal (EO revocation) is invisible to the model. In Q2, a legislative amendment changes a specific dollar figure that the model can only approximate from outdated data.
Chain-of-thought prompting amplifies these failures rather than mitigating them. By encouraging the model to reason step-by-step from stale premises, CoT produces internally coherent but factually incorrect justifications. This “confident confabulation” effect explains why CoT accuracy (30.0%) falls below even zero-shot (58.0%) on LegalSearchQA.
L-MARS resolves both cases by retrieving current authoritative sources (whitehouse.gov, IRS.gov) and grounding the answer in retrieved evidence rather than parametric memory. The system’s ability to access post-training information is the primary driver of its 38-percentage-point improvement over zero-shot on this benchmark.